# Foundation Models for Atomistic Simulation of Chemistry and Materials

### v1

@@ 3.3 Model-Agnostic Training Strategies
Post-training strategies. After pre-training a FM, post-training is an important procedure to further refine and optimize the model to a given downstream application. For domain specific tasks, such as energy and force prediction, modest amounts of higher-quality and labeled data are used within a standard supervised fine-tuning (SFT) context. Model distillation is another post-training approach for fine-tuning. In model distillation, a simpler model (the "student") is trained to replicate the behavior and performance of a larger, more complex model (the "teacher", or FM in this case). This process involves transferring knowledge from the teacher to the student by using the teacher’s predictions or internal representations as targets for training the student model.
Figure 3: Example of a post-training strategy for machine learning interatomic potentials via model distillation. One starts with an MLIP FM that has been trained on a large quantity of diverse data. A series of smaller MLIPs can then be trained via a knowledge distillation procedure, enabling specialization and speed for a specific task while retaining the general-purpose representations learned by the large-scale FM. Figure is reproduced from reference [133].
@@ 3.4 Criteria for a Successful Foundation Model
Even if the MLIP universality within a specific domain can be evaluated through OOD tests, the leaderboards do not seek to show that the resulting models are foundational, i.e. that they obey scaling laws or are derived from diverse data or pre-training strategies or that they exhibit superior performance on a broad range of down-stream tasks after fine-tuning. While the large size of the OC20 training set makes it a viable arena in which to evaluate model scalability, e.g. via performance with  1M vs  10M vs  100M training data, only performance on the full training dataset is featured on the leaderboard. Further, both OC20 and MatBench Discovery focus on well-behaved relaxations but without consideration of other emergent properties such as MD stability or PES smoothness, which has been shown to be independent of just test errors for energies and forces.[146, 154, 143, 155] However the recently released "NNP Arena" by Rowan Sci (https://benchmarks.rowansci.com/) is a step in the right direction, by evaluating energy error for both pre-trained MLIPs (and pure DFT models) against gold-standard coupled cluster reference values, without any concern for what data was used in MLIP training or how MLIP training was performed. NNP Arena further evaluates inference speed, which previous leaderboards haven’t considered but which may be a very important consideration for the future.
@@ 4 Discussion and Future Directions
The primary promise of a MLIP foundation model is that it would enable massively impactful investigation of systems ranging from metal-organic frameworks to enzymes to electrocatalysis beyond current physics-only models, while avoiding the need to build each ML model from the ground-up for any new chemical system of interest. One key insight that has been observed in other areas of machine learning is that methods that may work in “low data” regimes do not necessarily work as well in higher data regimes. As data size increases, as we have summarized for the many large-scale community and data efforts in chemistry and materials, can ML architectures scale accordingly to handle such data? Currently, the two paradigms are to use inductive biases to address physical constraints, often requiring less data[31, 32], with an alternative philosophy being that such physical laws can be easily learned when there is enough data, and when model capacity and compute is correspondingly scaled to take advantage of this data. Constrained MLIPs or learning energy and forces together ensures that conservative forces are available for the primary downstream MD task, but post-training strategies can take non-conservative MLIPs and distill them into conservative MLIPs, enabling both initial training at scale and physical soundness at deployment. Hence we posit here that incorporation of architectural inductive biases is not the only way to create sound MLIP FMs, a direction that is currently an underexplored paradigm in chemistry and materials. It raises the interesting question whether the chemical and materials sciences will escape the “bitter lesson”, or whether our science domains may also conclude that data and scalability is superior to built-in constraints, as has come to pass in other fields such as robotics and weather that are also grounded in physics. Early conclusions along these line are already evident in the pretraining model Uni-Mol2 that has characterized the scaling correlations of decreasing validation error with increasing model size, dataset size, and computational resources.[64]
@@
We note that the success of foundational LLMs and LVMs has critically depended on scraping vast collections of text and images from the internet, i.e. data obtained via implied or enforced open access. But in the context of DFT data for training MLIPs, this assumption of open access to data is not always adhered to when commercial players are involved. For instance, the datasets used to train Google’s GNoME model for crystal energies and Microsoft Research reported MatterSim have not been released for public use. In contrast, the FAIR Chemistry team at Meta, responsible for the largest MLIP training datasets built to-date, has consistently released their datasets (and trained models) fully open-source, a value to the scientific community that cannot be understated.
@@
Finally, while current MLIP testing platforms do provide a substantial resource to the community, there are still significant challenges and opportunities in the context of MLIP FMs. For example, there is not always a direct correlation between reduced force errors observed in test set performance and its actual performance in MD simulations.[146, 154, 143, 155]. Current MLIP leaderboards also do not yet explicitly evaluate model scalability with respect to the amount of training data or learnable parameters, nor do they seek to quantify the "foundational capacity" of the model via standardized fine-tuning across chemical and materials domains and properties. Scaling laws being the defining property of a FM, MLIPs should demonstrate continuous improvement with more parameters, more data, and more compute. Ultimately the FM practitioner must demonstrate the success of a FM via fine-tuned models that are fast and specialized - and which are more accurate than a model trained from scratch on data specific to that task/domain.
The development of a scalable FM will require using efficient algorithms to perform pre-training and fine-tuning. In particular, efficient mechanism for storing and retrieving data for batched training, dynamic load balancing methods that can accommodate variations in graph sizes, and communication hiding techniques that overlap computation and message passing will be important for achieving a rapid turnaround. A practical challenge is the computational resources required for the demonstration of foundational model capabilities at large scale. It will require world-class high performance computing beyond the academic lab scale and the first place to turn to are the leadership computing levels at the national labs. It may require an even larger federation of government agencies with industry, and international cooperation, modeled on the paradigm of scale of the Human Genome Project.
5 ACKNOWLEDGMENTS
The authors thank Eric Qu, Yingze Wang, Joseph Cavanagh, Kunyang O. Sun, Kareem Hegazy, Joseph P. Heindel, Peichen Zhong, Anup Kumar, Senwei Liang, Selim Sami, R. Allen LaCour, Sruthy K. Chandy, Ruoqi Zhao, and Dorian Bagni for their contributions in discussions on August 6, 2024 at UC Berkeley. T.H.-G. thanks the CPIMS program, Office of Science, Office of Basic Energy Sciences, Chemical Sciences Division of the U.S. Department of Energy under Contract DE-AC02-05CH11231 for support of the machine learning research. S.M.B. acknowledges support by the Center for High Precision Patterning Science (CHiPPS), an Energy Frontier Research Center funded by the U.S. Department of Energy, Office of Science, Basic Energy Sciences at Lawrence Berkeley National Laboratory under contract DE-AC02-05CH11231. This work was also supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing, and Office of Basic Energy Sciences, via the Scientific Discovery through Advanced Computing (SciDAC) program for the mathematical and computing foundations of machine learning models. This work was also supported by the U.S. Department of Energy, Office of Science, Energy Earthshot initiatives as part of the Center for Ionomer-based Water Electrolysis at Lawrence Berkeley National Laboratory under Award Number DE-AC02-05CH11231, as well as Toyota Research Institute as part of the Synthesis Advanced Research Challenge. This work used computational resources provided by the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility operated under Contract DE-AC02-05CH11231.
6 AUTHOR CONTRIBUTIONS STATEMENT
S.M.B., B.C., A.K., and T.H.-G. defined the topics and scope of the perspective. E.C.-Y.Y., Y.L., J.C., S.R., T.K., S.V., W.X. made the figures. E.C.-Y.Y., Y.L., J.C., S.R., T.K., S.V., P.Z., W.X., S.M.B., B.C., A.K., and T.H.-G. wrote the paper. All authors discussed the results and made comments and edits to the manuscript.
7 COMPETING INTERESTS STATEMENT
The authors declare no competing interests.

### v3

@@ 3.3 Model-Agnostic Training Strategies
Post-training strategies. After pre-training a FM, post-training is an important procedure to further refine and optimize the model to a given downstream application. For domain specific tasks, such as energy and force prediction, modest amounts of higher-quality and labeled data are used within a standard supervised fine-tuning (SFT) context. Model distillation is another post-training approach for fine-tuning. In model distillation, a simpler model (the "student") is trained to replicate the behavior and performance of a larger, more complex model (the "teacher", or FM in this case). This process involves transferring knowledge from the teacher to the student by using the teacher’s predictions or internal representations as targets for training the student model. The use of teacher-student constructs frequently allows to perform more computationally efficient training, particularly for multi-label training cases.
Figure 3: Example of a post-training strategy for machine learning interatomic potentials via model distillation. One starts with an MLIP FM that has been trained on a large quantity of diverse data. A series of smaller MLIPs can then be trained via a knowledge distillation procedure, enabling specialization and speed for a specific task while retaining the general-purpose representations learned by the large-scale FM. Reproduced with permission from Ref. [160].
@@ 3.4 Criteria for a Successful Foundation Model
Even if the MLIP universality within a specific domain can be evaluated through OOD tests, the leaderboards do not seek to show that the resulting models are foundational, i.e. that they obey scaling laws or are derived from diverse data or pre-training strategies or that they exhibit superior performance on a broad range of down-stream tasks after fine-tuning. While the large size of the OC20 training set makes it a viable arena in which to evaluate model scalability, e.g. via performance with  1M vs  10M vs  100M training data, only performance on the full training dataset is featured on the leaderboard. Further, both OC20 and MatBench Discovery focus on well-behaved relaxations but without consideration of other emergent properties such as MD stability or PES smoothness, which has been shown to be independent of just test errors for energies and forces.[173, 181, 170, 182] However the recently released NNP Arena[183] is a step in the right direction, by evaluating energy error for both pre-trained MLIPs (and pure DFT models) against gold-standard coupled cluster reference values, without any concern for what data was used in MLIP training or how MLIP training was performed. NNP Arena further evaluates inference speed, which previous leaderboards haven’t considered but which may be a very important consideration for the future. Another important direction pursued in the recent MLIP Arena[184] involves assessing the extent to which models comply with fundamental physics, including energy conservation, asymptotic behavior, smoothness, symmetry, and simulation stability. These metrics do not require additional labels, yet they carry substantial implications for the reliability and applicability of the models in practical settings.
3.5 Toward Foundation Models in Chemistry and Materials
At present all MLIPs are not true FMs by all criteria outlined in Section 2, but there are a few prototype MLIP FMs on the horizon that are proving more accurate, to more comprehensively cover the periodic table for molecular systems, and also show the ability to comply with heuristic scaling laws. In particular the pretraining models such as Efficiently Scaled Attention Interatomic Potential (EScAIP)[44] and Uni-Mol2[75] have characterized the scaling correlations of decreasing validation error with increasing model size, dataset size, and computational resources. This would suggest that emergent capabilities might follow, such as predicting energy labels from unlabeled data as shown for Uni-Mol2[75]. The fact that LLMs have demonstrated such scaling laws has led to emergent properties in chemistry for generative tasks in small molecule drugs[7] and their synthesis[9], although the LLMs were not trained on SMILES strings or chemical reaction templates.
Since the initial writing of this manuscript, two important contributions have been released which substantially advance the state-of-the-art for atomistic simulation: the Open Molecules 2025 (OMol25) dataset[80], and the Universal Model of Atoms pre-trained MLIP[185]. OMol25 is the largest ever molecular DFT dataset for training MLIPs, with over 100 million snapshots at an excellent level of DFT ($\omega$B97M-V/def2-TZVPD), where snapshots include up to 350 atoms, up to ten unpaired electrons, total charges ranging from -10e to +10e, and 83 elements of the periodic table. OMol25 broadly covers small molecules, electrolytes, biomolecules, and metal complexes, using more than ten distinct methods for structural sampling.[80] OMol25 thus makes substantial progress towards high-quality data coverage beyond organics, such as non-covalent interactions, spin and charge species, complex reactions, and long-range interactions, nearly half of the areas outlined in Figure 2. In addition to the $>$100 million training data, OMol25 also includes explicit out-of-distribution test sets and novel model evaluations, which will populate a public leaderboard in the near future, and baseline trained MLIPs.
Simultaneous with OMol25’s release was the release of UMA[185], which is trained on OMol25[80], OMat24[61], ODAC23[96], OC20[132], and the novel Open Molecular Crystal dataset (unpublished), totaling nearly 500 million training data - the largest and most diverse training set ever used to train a single MLIP. UMA also demonstrates scaling laws for both number of learnable parameters and compute, thus coming the closest to our vision of a foundation model for atomistic simulation. UMA is based on the eSEN architecture[186] and further employs a novel mixture of linear experts (MoLE) for two key benefits: 1) the MoLE causes a model trained on all datsets simultaneously, in a multi-task fashion (where each distinct training dataset is a distinct "task"), to exhibit superior performance to models trained separately on each dataset. 2) the MoLE allows a huge number of learnable parameters to be active during training, but a much smaller number to be active at inference time, allowing the model to effectively leverage the vast training dataset while still remaining fairly fast and memory efficient for subsequent simulations.
UMA and eSEN-OMol now top Rowan’s NNP Arena leaderboard[183] for neutral, closed-shell, organic molecular energies, with errors against coupled cluster for the GMTKN55 benchmark set falling by nearly 5 kcal/mol compared with the previous best model. Digging more deeply into the GMTKN55 test set[187], one can see that UMA’s energies are already lower error with respect to CCSD(T) than $\omega$B97M-D3BJ/def2-QZVP DFT for variable charge, variable spin, and metal-containing large systems and barrier heights, while being nearly comparable for intermolecular non-covalent interactions. However, the OMol25 evaluation tasks demonstrate that there is still substantial room for improvement on charge, spin, and long-range interactions, where error against reference DFT data remains very large ($>100$ meV) likely due to the naive handling of charge and spin and the 6Å graph cutoff, emphasizing the key areas where further MLIP architecture development is needed.
@@ 4 Discussion and Future Directions
The primary promise of a MLIP foundation model is that it would enable massively impactful investigation of systems ranging from metal-organic frameworks to enzymes to electrocatalysis beyond current physics-only models, while avoiding the need to build each ML model from the ground-up for any new chemical system of interest. One key insight that has been observed in other areas of machine learning is that methods that may work in “low data” regimes do not necessarily work as well in higher data regimes. As data size increases, as we have summarized for the many large-scale community and data efforts in chemistry and materials, can ML architectures scale accordingly to handle such data? Currently, the two paradigms are to use inductive biases to address physical constraints, often requiring less data[41, 42], with an alternative philosophy being that such physical laws can be easily learned when there is enough data, and when model capacity and compute is correspondingly scaled to take advantage of this data. Constrained MLIPs or learning energy and forces together ensures that conservative forces are available for the primary downstream MD task[159], but post-training strategies can take non-conservative MLIPs and distill them into conservative MLIPs[160], enabling both initial training at scale and physical soundness at deployment. Hence we posit here that incorporation of architectural inductive biases is not the only way to create sound MLIP FMs, a direction that is currently an underexplored paradigm in chemistry and materials. It raises the interesting question whether the chemical and materials sciences will escape the “bitter lesson”, or whether our science domains may also conclude that data and scalability is superior to built-in constraints, as has come to pass in other fields such as robotics and weather that are also grounded in physics.
@@
While current MLIP testing platforms do provide a substantial resource to the community, there are still significant challenges and opportunities in the context of MLIP FMs. For example, there is not always a direct correlation between reduced force errors observed in test set performance and its actual performance in MD simulations.[173, 181, 170, 182]. Current MLIP leaderboards also do not yet explicitly evaluate model scalability with respect to the amount of training data or learnable parameters, nor do they seek to quantify the "foundational capacity" of the model via standardized fine-tuning across chemical and materials domains and properties. Scaling laws being the defining property of a FM, MLIPs should demonstrate continuous improvement with more parameters, more data, and more compute. Ultimately the FM practitioner must demonstrate the success of a FM via fine-tuned models that are fast and specialized - and which are more accurate than a model trained from scratch on data specific to that task/domain. The development of a scalable FM will also require using efficient algorithms to perform pre-training and fine-tuning. In particular, efficient mechanism for storing and retrieving data for batched training, dynamic load balancing methods that can accommodate variations in graph sizes, and communication hiding techniques that overlap computation and message passing will be important for achieving a rapid turnaround.
We note that the success of foundational LLMs and LVMs has critically depended on acquisition of vast collections of text and images from the internet, i.e. data obtained via implied or enforced open access. But in the context of DFT data for training MLIPs, this assumption of open access to data is not always adhered to when commercial players are involved. For instance, the datasets used to train Google’s GNoME model for crystal energies and Microsoft Research reported MatterSim have not been released for public use. In contrast, the FAIR Chemistry team at Meta, responsible for the largest MLIP training datasets built to-date, has consistently released their datasets (and trained models) fully open-source, a value to the scientific community that cannot be understated. In a related vein, MLIP models have benefited tremendously from industry-released software tools such as TensorFlow[188], Keras[189], and PyTorch[190]. However, open-access development has not occurred for LLMs such as ChatGPT[191], and the loss of implementation details on how LLM FMs work could limit our ability to develop similarly powerful MLIP FMs. We note that while DeepSeek-V3 is open source, and can deliver high-quality results at a lower cost compared to close-source LLMs, it is either restricted or outright banned in several countries due to concerns about ethics, privacy, and security. As is needed for all emerging technologies, training in ethical development and use of AI/ML should remain a top priority in the physical sciences.[192] An additional practical challenge is the computational resources required for the demonstration of FM capabilities at large scale. It will require world-class high performance computing beyond the academic lab scale and the first place to turn to are the leadership computing levels at the national labs. The combined issues of open source data, software, models, politics, and ethics training will likely require an even larger federation of government agencies with industry, and international cooperation, modeled on the paradigm of scale of the Human Genome Project[193].